Let V and W be two vector spaces over the field F and T, T1, T2 be linear transformations from

V to W. Prove the following: (2+3=5 marks)

(a) rank(αT) = rank(T), for all α ∈ F and α 6= 0.

(b) |rank(T1) − rank(T2)| ≤ rank(T1 + T2) ≤ rank(T1) + rank(T2).


1
Expert's answer
2021-11-02T18:07:06-0400

Let V and W be vector spaces. A linear transformation T : V → W is a function from V to W such that T(u + v) = T(u) + T(v) and T(cu) = cT(u) for all u, v ∈ V and for all c ∈ R


a)

rank(T) is dimension of T(v).

"\alpha T(v)=T(\alpha v)"

dimension of "T(\alpha v)=" dimension of "T( v)"

So, rank(αT) = rank(T)


b)

since image of (T1+T2) is subspace of (image of T1+image of T2):

"rank(T1 + T2) \u2264 rank(T1) + rank(T2)"


since image of T1 and image of T2 are subspaces of image of (T1+T2):

"|rank(T1) \u2212 rank(T2)| \u2264 rank(T1 + T2)"


so,

"|rank(T1) \u2212 rank(T2)| \u2264 rank(T1 + T2) \u2264 rank(T1) + rank(T2)"